A flow of 834 cfs is conveyed in a 10-ft-wide rectangular channel at a depth of 6.0 ft. What is the specific energy of the channel? Is this flow subcritical or supercritical? If n = 0.014, what channel slope must be provided to produce a uniform flow at this depth? Check your solution using appropriate computer software.


1
Expert's answer
2022-01-20T13:50:43-0500

Solution

Given data are


Flow rate Q=834cfs=23.35m3/sec.

Width w=10ft=3.048m

n=0.0014

Depth d=6ft=1.828m

Now

Specific energy

"E=\frac{Q^2}{2gA^2}+d"

Putting all values

"E=\frac{(23.35)^2}{2\times 9.8\times (3.048\times1.828)^2}+1.828"

E=2.723J

Now

For checking the flow rate

"F_r=\frac{\frac{Q}{A}}{\sqrt{gd}}"

"F_r=\frac{\frac{23.35}{3.048\times1.828}}{\sqrt{9.8\times1.828}}"

So "F_r=0.889<1"

So this is subcritical.

So

Slope for uniform flow

S="\frac{1.49AR^{0.66}D^{0.5}}{n}"

Putting all values we get

S="1.298\times10^{-3}"